A positive and a negative charge, each of magnitude 2.6 x 10–5 C, are separated by a distance of 25 cm. Find the force on each of the particles

1
Expert's answer
2021-10-19T15:09:48-0400

Given:

"q=2.6*10^{-5}\:\rm C"

"r=0.25\:\rm m"

"k=9*10^9\:\rm N\cdot m^2/C^2"


The Coulomb law says

"F=k\frac{q_1q_2}{r^2}=k\frac{q^2}{r^2}"

"F=9*10^9*\frac{(2.6*10^{-5})^2}{0.25^2}=97\:\rm N"

The force is attraction.